OpenStudy (anonymous):

If 62.8 g Ca react with 92.3 g HCl according to the reaction below, how many grams of hydrogen gas will be produced, and how many grams of the excess reactant will be left over? Unbalanced equation: Ca + HCl → CaCl2 + H2 Be sure to show all of your work. @Nurali

OpenStudy (somy):

balance the reaction first

OpenStudy (anonymous):

I don't know how, I am so lost in my chemistry class it's scary

OpenStudy (nurali):

Balance the equation Ca + 2HCl -------> CaCl2 + H2 moles = mass / molar mass moles Ca = 62.8 g / 40.08 g/mol = 1.5669 mol moles HCl = 92.3 g / 36.458 g/mol = 2.532 mol The balanced equation tells you that 1 mole Ca needs 2 moles HCl to fully react So moles HCl needed to react all the Ca = 2 x moles Ca = 2 x 1.5669 mol = 3.1338 mol HCl needed But you are only give 2.532 mol HCl, which means you don't have enough to react all the Ca, so HCl is limiting and Ca is in excess. The maximum amount of H2 possible is if all the limiting reagent reacts. 2 moles HCl reacts to give 1 mole H2 So moles H2 possible = 1/2 x moles HCl = 1.2 x 2.532 mol = 1.266 moles H2 mass H2 = molar mass x moles = 2.016 g/mol x 1.266 mol = 2.55 g H2 work out how many moles of Ca you will need 1 mole Ca reacts with 2 moles HCl So moles Ca = 1/2 x moles HCl = 1.266 mol Ca used moles Ca left over = amount provided - amount used = 1.5669 mol - 1.266 mol = 0.3009 mol mass = molar mass x moles = 40.08 g/mol x 0.3009 mol = 12.1 g

OpenStudy (somy):

Ca + HCl → CaCl2 + H2 count number of Ca, H, and Cl on the right and left side of the arrow left side: Ca - 1 H - 1 Cl - 1 right side Ca -1 H -2 Cl -2 as you can see there are two of H and Cl on the right side so we can write 2HCl on the left side to equal their numbers on both sides Ca + 2HCl → CaCl2 + H2

OpenStudy (somy):

wow that was fast? o.o @Nurali

OpenStudy (anonymous):

Wow. How do y'all know this stuff? Somy, can you give a medal to Nurali so I can give you one?

OpenStudy (nurali):

Thanks @Somy

OpenStudy (somy):

i already did lol that's too good

OpenStudy (anonymous):

Thank you guys!!!

OpenStudy (somy):

no problem tho credits go mostly to @Nurali :3

OpenStudy (nurali):

My Pleasure @Redneck1444